30cm of hydrogen diffuses out in 2 seconds. How long will it take 90cm of oxygen to diffuse out under the same conditions
1
Expert's answer
2016-02-12T05:34:01-0500
According to Graham's law:
rateH22 x MH2 = rateO22 x MO2
M (H2)=2 g/mol
M (O2) = 32 g/mol
RateH2: 30/2=15 cm/s.
RateO2=(rateH22 x MH2 )/MO2
RateO2=(152 x 2 )/32 = 14.1 cm/s
So that 90 cm of oxygen under the same conditions will be diffused during:
T=90/14.1=6.4 s